# The Intrinsic Flattening of Extragalactic Stellar Disks

### v1

@@ 1 Introduction
Our paper is organised as follows. Our brief historical review in Sec. 2 reveals that most previous measurements of $c/a$ have largely relied on ground-based blue sensitive, and thus dust-extinction prone, material and lacked a precise definition of the $c$ and $a$. We suggest a method based on infrared digital surveys of disk galaxies. In Sec. 3, we present the data used in our analysis, ranging from science-ready images to disk decompositions from other works. Our data reduction methodology to identify isophotes that characterise the disk’s structure, as well as relevant statistical concerns, are addressed in Sec. 4. Results are presented in Sec. 5. The shape of the intrinsic flattening distribution, and whether it is normal, is considered along with correlations between our (model-independent) $c/a$ values and structural galaxy properties, including similar measures from parametric (model-dependent) disk decompositions. We also test for any biases due to galactic environments. A discussion of our main findings is presented in Sec. 6, including a discussion between the connection of model independent measures of $c$ versus model dependent estimates of a disk scale-height. Future avenues are discussed in Sec. 7 and followed by a summary in Sec. 8. App. A discusses potential resolution effects on the measurement of $c/a$.
@@ 2 Historical Remarks
The definition of an unbiased estimate of the “true” flattening of disk systems is challenging since disks are not monolithic slabs; within the optical radius, the vertical height of a disk decreases (often exponentially) with galactocentric radius (e.g., van der Kruit & Searle, 1981; de Grijs et al., 1997; de Grijs, 1998; van der Kruit & Freeman, 2011; Bizyaev et al., 2014), and may increase again in the outskirts due to disk flaring (Olling, 1995; Ossa-Fuentes et al., 2023). Of issue then is the availability and widespread application of a rigorous definition of $a$ and $c$ which appropriately describes the disk.
@@ 3.1 The Spitzer Survey of Stellar Structures in Galaxies Sample
S4G galaxies were classified according to the CVRHS system by Buta et al. (2015), with updated $T$-type classifications (Watkins et al., 2022). The morphological distribution of galaxies of each $T$-type (and corresponding morphology) in our sample is presented in Table 1. It should noted that galaxy classification of edge-on systems is especially problematic, though the reduced dust extinction in the IRAC bandpass is a significant improvement for this task over bluer dust-proned images. Separating barred from unbarred systems for edge-on systems is also challenging, unless a clear peanut bulge can be seen. We address barred systems in Sec. 5.1 but caution about hasty interpretations.
@@ 3.2 Scale Parameters
As described below (Sec. 4.1), our non-parametric method to estimate the intrinsic flattening, $c/a$, of galaxy disks relies on isophotal fitting of (transparent) galaxy images. Galaxy surface brightness (SB) profiles, obtained from radial cuts for edge-on systems, can also be used to measure relevant structural parameters such as the disk scale height, $z_{0}$, and scale length, $h$, and isolate substructure. We will compare in Sec. 5 our non-parametric method of measuring $c/a$ to the parametric ratio of $z_{0}$ and $h$, the disk scale height and length. To this end, we have used Comerón et al. (2018)’s decompositions of (major axis) SB profiles for our S4G galaxies to source values of the scale heights for the thin and thick disks. We have also extracted disk scale lengths from the (major axis) SB profiles of our S4G galaxies by fitting a single exponential to the disk-dominated regions of the profiles.
@@ 3.3 Global Parameters
Comerón et al. (2018) derived $M_{\text{CMC}}$, of which we avail ourselves, from their SB profile decomposition of S4G galaxies. Circular velocity ($V_{\text{rot}}$) data were also taken from the compilation of Comerón et al. (2018). If possible, Comerón et al. (2018) sourced data for the derivation of $V_{\text{rot}}$ from the Extragalactic Distance Database (EDD; Tully et al. 2009), which provides information and analysis of H i l̇ine widths to determine $V_{\text{rot}}$. Otherwise, Comerón et al. (2018) sourced $V_{\text{rot}}$ data from stellar absorption line studies (see Comerón et al. 2018 for a comprehensive list of sources).
@@ 4.1 Characteristic Cuts
Our measurement of the intrinsic flattening of the stellar disk component used two ensembles of isophotes that span large sections of the disk. The constituent isophotes are those that intersect the galactic midplane within two bins defined by the boundaries $0.2\,R_{25}\leq|x|\leq 0.5\,R_{25}$, for the inner cut, and $0.5\,R_{25}\leq|x|\leq 0.8\,R_{25}$, for the outer cut. These are shown in Fig. 2. While this choice of cuts is somewhat arbitrary, they match those of Comerón et al. (2018), allowing for direct comparisons. These cuts also roughly divide isophotes dominated by the thin and thick disk light, respectively.
@@
Figure 2: Top: Radial profile of intrinsic flattening for IC 0335 (as a typical example). The teal upward hatched area the red downward hatched area define the inner cut from $0.2\!-\!0.5\,a/R_{25}$ and the outer cut over $0.5\!-\!0.8\,a/R_{25}$, respectively. Bottom: Stacked density distribution of all 133 intrinsic flattening radial profiles in our sample. The horizontal black line marks $c/a=0.12$. See text for details.
@@ 5.1 The Distribution of Intrinsic Flattening vs Morphology
For lenticular galaxies ($T$ = [$-3$, 0]), we find that earlier systems appear to be thicker, albeit with large scatter. The thickening towards earlier types can be attributed to the evolutionary processes that shaped them; galaxy interactions increase the velocity dispersion in the disk while gas depletion halts the formation of new, low velocity dispersion stars (Comerón et al., 2016; Pinna et al., 2019a, b). Our result above contrasts with HG84’s constant intrinsic flattening ($c/a=0.23$) for all lenticulars.
@@ 5.3 Correlations between c/a and Model-Dependent Parameters
To investigate these trends, we have computed the average value of $a$ for both inner and outer bins as $a_{\rm{in}}=0.35\,R_{25}$ and $a_{\rm{out}}=0.65\,R_{25}$, respectively. To estimate the error on these values, we assigned both values of $a$ the maximum possible uncertainty of $a_{\rm{err}}=0.15\,R_{25}$, which is the distance from the center to the edge of either cut. As a result of this choice, the percent errors on $a$ will be identical for a given cut: $42\%$ error for the inner cut and $23\%$ error for the outer cut. It follows that the semi-minor axis of the characteristic cut is $c_{i}=a_{i}(c/a)_{i}\mathrm{,}$ where $i$ denotes either the inner or outer cut. The uncertainty on $c/a$ and $a$ is propagated through to $c$. With values of $a$ in place, values of $c$ emerge directly from the ratio $c/a$.
@@ 6.2 The Effectiveness of the Method of Characteristic Cuts
To assess whether the method of characteristic cuts produces most accurate measurements of intrinsic flattening, we compared our results to those derived from the more straightforward method of choosing a singular isophote. A traditional isophotal reference is $\mu_{B}=25$ mag arcsec-2 . Fig. 11 shows the error-weighted distribution of our sample spiral and lenticular’s intrinsic flattening measured from the isophote of radius $r=R_{25}$, or whichever isophote whose radius was closest to $R_{25}$.
@@ 8 Summary
Defining the size of a galaxy based on its projected image on the sky is especially challenging and requires care. We have proposed the method of characteristic cuts for measuring the intrinsic flattening of disk galaxies and applied it to a sample of 141 (nearly) edge-on galaxies from the S4G imaged in $3.6\ \mu$m. We used AutoProf, a Python-based automated tool for the photometry of astronomical images (Stone et al., 2021), to perform isophotal fits on the $3.6\ \mu$m images of each galaxy. With these isophotes, we established the characteristic cuts of each galaxy by taking the average axial ratio $c/a$ of isophotes falling within axial bins of $0.2\,R_{25}\leq|x|\leq 0.5\,R_{25}$ and $0.5\,R_{25}\leq|x|\leq 0.8\,R_{25}$, which we referred to as the inner and outer cuts, respectively. Those galaxies which showed clear visual indications of poor fits or had large $c/a$ variances within either cut were rejected, leaving us with a sample of 133 galaxies with good fits. We also retrieved surface brightness profiles and thin and thick disks decompostions from Comerón et al. (2018) in addition to various physical parameters from Muñoz-Mateos et al. (2015). Our findings are summarized here:
@@ 9 Acknowledgments
SC(Canada) and CS acknowledge generous support from the Natural Sciences and Engineering Research Council of Canada and Queen’s University through various scholarships and grants. CS also acknowledges support from the Canadian Institute for Theoretical Astrophysics (CITA) National Fellowship program. SC(Spain) also acknowledges funding from the State Research Agency (AEI-MCINN) of the Spanish Ministry of Science and Innovation under the grant “Thick discs, relics of the infancy of galaxies” with reference PID2020-113213GA-I00. This research project was initiated under the watchful eye of Simon Díaz García. We are grateful for his precious suggestions and contributions. Peter Erwin is thanked for insightful discussions and references.

### v3

@@ 1 Introduction
Our paper is organised as follows. Our brief historical review in Sec. 2 reveals that most previous measurements of $c/a$ have largely relied on ground-based blue sensitive, and thus dust-extinction prone, material and lacked a precise definition of the $c$ and $a$. We suggest a method based on infrared digital surveys of disk galaxies. In Sec. 3, we present the data used in our analysis, ranging from science-ready images to disk decompositions from other works. Our data reduction methodology to identify isophotes that characterise the disk’s structure, as well as relevant statistical concerns, are addressed in Sec. 4. Results are presented in Sec. 5. The shape of the intrinsic flattening distribution, and whether it is normal, is considered along with correlations between our (model-independent) $c/a$ values and structural galaxy properties, including similar measures from parametric (model-dependent) disk decompositions. We also test for any biases due to galactic environments. A discussion of our main findings is presented in Sec. 6, including a discussion contrasting model-independent and model-dependent estimates of intrinsic flattening. Future avenues are discussed in Sec. 7 and followed by a summary in Sec. 8. App. A discusses potential resolution effects on the measurement of $c/a$.
@@ 2 Historical Remarks
The definition of an unbiased estimate of the “true” flattening of disk systems is challenging since disks are not monolithic slabs; within the optical radius, the vertical height of a disk decreases (often exponentially) with galactocentric radius (e.g., van der Kruit & Searle, 1981; de Grijs et al., 1997; de Grijs, 1998; van der Kruit & Freeman, 2011; Bizyaev et al., 2014), and may increase again in the outskirts due to disk flaring or warping (Olling, 1995; Ossa-Fuentes et al., 2023). Of issue then is the availability and widespread application of a rigorous definition of $a$ and $c$ which appropriately describes the disk.
@@ 3.1 The Spitzer Survey of Stellar Structures in Galaxies Sample
S4G galaxies were classified according to the CVRHS system by Buta et al. (2015), with updated $T$-type classifications (Watkins et al., 2022). The morphological distribution of galaxies of each $T$-type (and corresponding morphology) in our sample is presented in Table 1. It should be noted that galaxy classification of edge-on systems is especially problematic, though the reduced dust extinction in the IRAC bandpass is a significant improvement for this task over bluer dust-proned images. Identifying barred systems in edge-on cases is also challenging, unless a peanut bulge is clearly seen. However, as seen in Sec. 5.1, barredness does not play a role in our assessment of $c/a$.
@@ 3.2 Scale Parameters
As described below (Sec. 4.1), our non-parametric method to estimate the intrinsic flattening, $c/a$, of galaxy disks relies on isophotal fitting of (transparent) galaxy images. Galaxy surface brightness (SB) profiles, obtained from radial cuts for edge-on systems, can also be used to measure relevant structural parameters such as the disk scale height, $z_{0}$, and scale length, $h$, and to isolate substructure. We will compare in Sec. 5 our non-parametric method of measuring $c/a$ to the parametric ratio of $z_{0}$ and $h$, the disk scale height and length. To this end, we have used Comerón et al. (2018)’s decompositions of (major axis) SB profiles for our S4G galaxies to source values of the scale heights for the thin and thick disks. We have also extracted disk scale lengths from the (major axis) SB profiles of our S4G galaxies by fitting a single exponential to the disk-dominated regions of the profiles.
@@ 3.3 Global Parameters
Comerón et al. (2018) derived $M_{\text{CMC}}$, of which we avail ourselves, from their SB profile decomposition of S4G galaxies. Circular velocity ($V_{\text{rot}}$) data were also taken from the compilation of Comerón et al. (2018). If possible, Comerón et al. (2018) sourced data for the derivation of $V_{\text{rot}}$ from the Extragalactic Distance Database (EDD; Tully et al. 2009), which provides information and analysis of H i l̇ine widths to determine $V_{\text{rot}}$. Otherwise, Comerón et al. (2018) sourced $V_{\text{rot}}$ data from stellar absorption line studies (see Comerón et al. 2018 for a comprehensive list of sources). Fig. 1 shows a comparison of our sub-sample (red histograms) with the full S4G sample. Our sub-sample favors slightly more massive, though H i  poorer, systems.
@@
Figure 1: Comparison of our sub-sample (red histograms) with the S4G catalogue (blue histograms) and (when available) the S4G early type galaxy catalogue (hatched histograms).
@@ 4.1 Characteristic Cuts
Our measurement of the intrinsic flattening of the stellar disk component used two ensembles of isophotes that span large sections of the disk. The constituent isophotes are those that intersect the galactic midplane within two cuts defined relative to the radius of the $25^{\rm{th}}$ B-band magnitude isophote, ${r_{25}}$. These cuts, depicted in Fig. 3, are defined by $0.2\,{r_{25}}\leq|a|\leq 0.5\,{r_{25}}$, for the inner cut, and $0.5\,{r_{25}}\leq|a|\leq 0.8\,{r_{25}}$, for the outer cut where $a$ is the galactocentric radius along the major axis of the galaxy. While this choice is somewhat arbitrary, they match those of Comerón et al. (2018), allowing for direct comparisons. These cuts also roughly divide isophotes dominated by the thin and thick disk light, respectively.
@@
Figure 3: Top: Radial profile of intrinsic flattening for IC 0335 (as a typical example). The teal upward hatched area and the red downward hatched area define the inner cut from $0.2\!-\!0.5\,a/{r_{25}}$ and the outer cut over $0.5\!-\!0.8\,a/{r_{25}}$, respectively. Bottom: Stacked density distribution of all 133 intrinsic flattening radial profiles in our sample. The horizontal black line marks $c/a=0.12$. See text for details.
@@ 5.1 The Distribution of Intrinsic Flattening vs Morphology
For lenticular galaxies ($T$ = [$-3$, 0]), we find that earlier systems appear to be thicker, albeit with large scatter, in agreement with older findings (Binney & de Vaucouleurs, 1981). The thickening towards earlier types can be attributed to the evolutionary processes that shaped them; galaxy interactions increase the velocity dispersion in the disk while gas depletion halts the formation of new, low velocity dispersion stars (Comerón et al., 2016; Pinna et al., 2019a, b). Our result above contrasts with HG84’s constant intrinsic flattening ($c/a=0.23$) for all lenticulars.
@@ 5.3 Correlations between c/a and Model-Dependent Parameters
To investigate these trends, we have computed the average value of $a$ for both inner and outer cuts as $a_{\rm{in}}=0.35\,{r_{25}}$ and $a_{\rm{out}}=0.65\,{r_{25}}$, respectively. To estimate the error on these values, we assigned both values of $a$ the maximum possible uncertainty of $a_{\rm{err}}=0.15\,{r_{25}}$, which is the distance from the center to the edge of either cut. As a result of this choice, the percent errors on $a$ will be identical for a given cut: $42\%$ error for the inner cut and $23\%$ error for the outer cut. It follows that the semi-minor axis of the characteristic cut is $c_{i}=a_{i}(c/a)_{i}\mathrm{,}$ where $i$ denotes either the inner or outer cut. The uncertainty on $c/a$ and $a$ is propagated through to $c$. With values of $a$ in place, values of $c$ emerge directly from the ratio $c/a$.
@@ 6.2 The Effectiveness of the Method of Characteristic Cuts
To assess whether the method of characteristic cuts produces the most accurate measurements of intrinsic flattening, we compared our results to those derived from the more straightforward method of choosing a singular isophote. A traditional isophotal reference is $\mu_{B}=25$ mag arcsec-2 . Fig. 12 shows the error-weighted distribution of our sample spiral and lenticular’s intrinsic flattening measured from the isophote of radius $r={r_{25}}$, or whichever isophote whose radius was closest to ${r_{25}}$.
@@ 8 Summary
Defining the size of a galaxy based on its projected image on the sky is especially challenging and requires care. We have proposed the method of characteristic cuts for measuring the intrinsic flattening of disk galaxies and applied it to a sample of 141 (nearly) edge-on galaxies from the S4G imaged in $3.6\ \mu$m. We used AutoProf, a Python-based automated tool for the photometry of astronomical images (Stone et al., 2021), to perform isophotal fits on the $3.6\ \mu$m images of each galaxy. With these isophotes, we established the characteristic cuts of each galaxy by taking the average axial ratio $c/a$ of isophotes falling within axial cuts of $0.2\,{r_{25}}\leq|a|\leq 0.5\,{r_{25}}$ and $0.5\,{r_{25}}\leq|a|\leq 0.8\,{r_{25}}$, which we referred to as the inner and outer cuts, respectively. Those galaxies which showed clear visual indications of poor fits or had large $c/a$ variances within either cut were rejected, leaving us with a sample of 133 galaxies with good fits. We also retrieved surface brightness profiles and thin and thick disks decompostions from Comerón et al. (2018) in addition to various physical parameters from Muñoz-Mateos et al. (2015). Our findings are summarized here:
@@ 9 Acknowledgments
This research project was initiated under the watchful eye of Simon Díaz García. We are grateful for his precious suggestions and contributions. Peter Erwin is also thanked for insightful discussions and references. SC(Canada) and CS acknowledge generous support from the Natural Sciences and Engineering Research Council of Canada and Queen’s University through various scholarships and grants. CS also acknowledges support from the Canadian Institute for Theoretical Astrophysics (CITA) National Fellowship program. SC(Spain) also acknowledges funding from the State Research Agency (AEI-MCINN) of the Spanish Ministry of Science and Innovation under the grant “Thick discs, relics of the infancy of galaxies” with reference PID2020-113213GA-I00.